OpenStudy (anonymous):

physical science : Near the surface of the Earth, the acceleration due to gravity is 9.8 m/sec/sec. After falling 6 sec, an object would have a velocity of a. 3.8 m/sec. b. 9.8 m/sec. c. 15.8 m/sec. d. 58.8 m/sec.

OpenStudy (anonymous):

58.8 m/sec

OpenStudy (anonymous):

Assuming you're ignoring air resistance, the equation v = u + at can be used: v = 0 + 9.8 * 6 = 58.8 m/s.

OpenStudy (anonymous):

thanks !!!!!!!!!!!!!:D

OpenStudy (anonymous):

(d) 58.8m/sec